OpenStudy (anonymous):

A solid gold cube with edge of 2" is melted down and recast as a cone with a 3" radius base. Find the height of the cone. A.) 8L(9pie) in. B.) 2piein. C.) 8L(3pie) in.

OpenStudy (anonymous):

cube volume is r^3 so 2^3 = 8 now this is melted to cone whose volume is 1/3(pi r^2 h) = 3 pi h so, 8 = 3 pi h

OpenStudy (anonymous):

Can u find h from here

OpenStudy (anonymous):

I have 1/3 (pi 8^2 h) = 3pih then I have 21.3pih =3pih so do I divide 3 from 21.3?

OpenStudy (anonymous):

For cone radius is 3 not 8, 8 is the volume of cube

OpenStudy (anonymous):

oh god, did I do my factoring right though, like how I solved it? Let me plug in a three.

OpenStudy (anonymous):

okay, this is what I got :

OpenStudy (anonymous):

I/3(pi3^2h) = 3pih then I times 1/3 with 9 to get 3(pih) =3pih so would it just be one when I divide it all out?

OpenStudy (anonymous):

equate 3pi h to 8 as volume of both cube & cylinder are equal 8 = 3 pi h h = 8/3 pi

OpenStudy (anonymous):

okay, now I get you! and since it can't be factored out I leave it as 8/3pi? :D

OpenStudy (anonymous):

yes thats the answer

OpenStudy (anonymous):

thank you so much! this really helped a lot :) <3 here is a barfing gnome for your help! :D

OpenStudy (anonymous):

thanks& yw:) always

OpenStudy (anonymous):

You are a bad retrice thanks again! :) <3

OpenStudy (anonymous):

*bad a$$*